Q12) A 19-gram bullet moving at 1689 m/s plunges into 2 kg of paraffin wax. The wax was initially at 25°C. Assuming that all the bullet's energy heats the wax, what is its final temperature (in ºC)? Take the mechanical equivalent of heat to be 4 J/cal and the specific heat of wax to be 0.7 cal/g °C.
Answer #1

I hope velocity is 168.9 m/sec

Kinetic energy of bullet = 0.5mv2

m is mass of bullet and v is speed if bullet

So KE = 0.5*(19*10-3)168.92 = 271 joules

this kinetic energy will convert into heat = Q

as Q = mc\DeltaT

m is mass of wax = 2 kg

c = specific heat of wax = 0.7*4 j/kg.K = 2.8 j/kg.K

\DeltaT is temperature difference

so 271 = 2*2.8*(T - 25)

T - 25 = 48.4 C

T = 73.4 C